Admission to a local aquarium is $10 for adults and $7 for children. The aquarium collected $6,459 from the sale of 783 tickets. How many children's tickets were sold?
Answer by TimothyLamb(4379)

x = number of adult tickets
y = number of child tickets
---
x + y = 783
10x + 7y = 6459
---
put the system of linear equations into standard form
---
x + y = 783
10x + 7y = 6459
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = number of adult tickets = 326
y = number of child tickets = 457
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---